# Street Gaussians without 3D Object Tracker

### v1

@@ Abstract
Realistic scene reconstruction in driving scenarios poses significant challenges due to fast-moving objects. Most existing methods rely on labor-intensive manual labeling of object poses to reconstruct dynamic objects in canonical space and move them based on these poses during rendering. While some approaches attempt to use 3D object trackers to replace manual annotations, the limited generalization of 3D trackers – caused by the scarcity of large-scale 3D datasets – results in inferior reconstructions in real-world settings. In contrast, 2D foundation models demonstrate strong generalization capabilities. To eliminate the reliance on 3D trackers and enhance robustness across diverse environments, we propose a stable object tracking module by leveraging associations from 2D deep trackers within a 3D object fusion strategy. We address inevitable tracking errors by further introducing a motion learning strategy in an implicit feature space that autonomously corrects trajectory errors and recovers missed detections. Experimental results on Waymo-NOTR datasets show we achieve state-of-the-art performance. Our code will be made publicly available.
Figure 1: Comparison of 3D tracker-based Street Gaussians [76] (left) and our approach (right). Existing methods heavily rely on object poses, but 3D trackers struggle with limited generalization, leading to flaws in novel view synthesis. In contrast, 2D foundation models show better generalization. Our approach leverages a 2D foundation model [72] for object tracking and learns point motion within an implicit feature space to autonomously correct tracking errors, improving robustness across diverse scenes.
@@ 1 Introduction
Although 2D tracking models demonstrate stronger generalization ability, tracking errors still can exist, especially under adverse conditions such as severe occlusions or distant objects. Since moving the object points solely based on the tracked object pose directly exposes any tracking errors, we aim to go beyond a straightforward reliance on tracked trajectories. Instead, we propose to learn point motion from the predicted trajectory, equipping the model to autonomously identify and correct tracking inaccuracies, recover missed detections, and infer motion in new time steps. An implicit representation is essential for this purpose: 1) It enables the model to refine trajectories without being constrained by bounding box tracks, facilitating smoother and more continuous motion that can adaptively respond to changes. 2) An implicit feature space offers versatility to move each point of an object in a different way, recognizing that vehicles are not strictly rigid objects — for example, doors can open or close. This approach makes it possible to capture subtle, dynamic changes within objects, ultimately producing more accurate reconstructions and enhancing the robustness of novel view synthesis in challenging scenarios. To this end, we leverage HexPlane representation [7] following 4DGS [69] and apply direct supervision on learned point motion using the predicted trajectory. However, applying 4DGS alone without the object tracking module and explicit motion supervision yields suboptimal results due to two main challenges: 1) Reconstructing fast-moving objects without tracking is difficult since aggregating points from different time frames results in a fragmented trail rather than a smooth and coherent representation. 2) Modeling rapid vehicle motion without explicit supervision proves challenging. For example, the projection of a Gaussian may fall outside the actual object area when the initial position is distant from its current position, impeding optimization since gradients cannot propagate from the object area to the Gaussian. These issues hinder precise motion modeling, as seen in S3Gaussian [29], which uses 4DGS without explicit supervision and thus struggles with moving cars, as demonstrated in our experiments.
@@
  • We achieve state-of-the-art performance on Waymo-NOTR dataset without relying on ground truth annotations.
@@ 3.3 Tracking Cars in 3D with Robust 2D Tracker
$\tau_{overlap}=\mathcal{Q}\cap\mathbf{T}_{t}^{-1}\mathcal{O}_{t},\quad\mathbf{T}_{t}=\operatorname{ICP}(\mathcal{O}_{t},\mathcal{O}).$ (4)
@@ 3.4 Learning Point Motion
Table 1: Comparison with state-of-the-art methods on NOTR dataset. StreetGS represents Street Gaussian [76]. The best and the second best results are denoted by pink and blue.
@@ 4.1 Experimental Setup
Datasets. The NeRF On-The-Road (NOTR) dataset, introduced by EmerNeRF [77], is a subset of the Waymo Open dataset [65] and includes various challenging driving scenarios, such as high-speed, exposure mismatch, dawn/dusk, gloomy, rainy, and night scenes. In contrast, many public LiDAR datasets such as nuScenes [6] and KITTI [44] mainly features simpler scenes with limited dynamic content. We leverage NOTR dynamic32 subsets, comprising 32 dynamic scenes to ensure a diverse benchmark for evaluating dynamic reconstruction.
Implementation Details. We use the Adam optimizer [35] with the same learning rate schedule as in 3DGS [33]. For long-sequence scene reconstruction, we follow S3Gaussian [29] and segment the scene into multiple clips. The multi-resolution HexPlane encoder has a base resolution of 64, upsampled by factors of 2 and 4 as in [69], and other hyperparameters match those in 3DGS [33]. As LiDAR points lack data for the sky region, we add a plane of points above the scene’s maximum height to represent it following [34]. For the Waymo-NOTR dataset, we strictly follow EmerNeRF’s experimental setup [77]. For camera images, we use three frontal cameras and resize them to a resolution of $960\times 640$ for both training and evaluation. For scene reconstruction, all samples within a log are utilized for training, probing the upper bound of each method’s performance. For novel view synthesis, we exclude every 10th time step, creating a 10% set of novel temporal views for evaluation. All experiments are conducted on a single NVIDIA RTX 3090 GPU, with training taking approximately 2 hours per video clip. The inference speed is around 80 FPS in resolution $960\times 640$.
Baseline Methods. We evaluate our approach against state-of-the-art methods, covering both NeRF-based and Gaussian-based 3D models. The NeRF-based methods include MARS [73] and EmerNeRF [77]. Gaussian-based methods include 3DGS [33], Street Gaussians [76] and S3Gaussian [29]. To ensure fair comparisons, we apply LiDAR point cloud initialization and add depth regularization to 3DGS. We borrow results of MARS [73] and EmerNeRF [77] from S3Gaussian [29]. Since Street Gaussians requires 3D tracking results as input, we evaluate it using detection results from VoxelNext [11] and associate the detections with SimpleTrack [56]. As NOTR is part of the training set for the perception task in the Waymo dataset, we use VoxelNext pretrained on the nuScenes dataset [6]. This setup also demonstrates the generalization capabilities of methods relying on 3D trackers.
@@ 4.2 Comparisons with the State-of-the-art
4.3 Initial Reconstruction Quality Evaluation
Figure 4: Histogram of reconstruction errors of our method and the 3D tracker-based reconstruction. All results with errors greater than 0.6 $m$ are placed in the rightmost bin.
To further evaluate the reconstruction quality of our object tracking module introduced in Sec. 3.3, we conduct an experiment on NOTR dataset. We use the ground truth 3D bounding box to crop the LiDAR point cloud within the box and transform it into the canonical space as the reference reconstruction. Only moving objects that are visible in at least two frames and contain more than 1,000 points in total are preserved. To associate ground truth objects with our predictions, we match each ground truth object with the nearest predicted object across all frames. Since our reconstruction is not in the same canonical space as the reference, we do the alignment before evaluation. For comparison, we repeat the same reconstruction process with 3D bounding boxes predicted by VoxelNext [11] as our baseline. We use Chamfer Distance to evaluate the reconstruction error.
The results are displayed as a histogram in Fig. 4. Note that all results with an error larger than 0.6 meters are placed in the rightmost bin. It can be seen that our method achieves lower reconstruction errors compared to the 3D tracker-based reconstruction. Specifically, 28.3% of the data points in the 3D tracker-based reconstruction have errors greater than 0.5 meters, while all data points in our method have errors below 0.5 meters, highlighting the robustness of our method.
@@ 4.4 Ablation Studies
No. Method PSNR↑ SSIM↑ DPSNR↑ DSSIM↑
@@
D 3D tracker 27.55 0.829 24.85 0.747
E GT pose 28.98 0.873 25.71 0.787
F Ours 28.85 0.867 25.58 0.779
@@
Effect of object tracking module. In Tab. 2 (A), we model dynamic points using 4DGS [69] without performing any form of tracking, and observe significant drops across all metrics. In Tab. 2 (B), we use the reconstruction of the tracking module for initialization but remove the vehicle motion loss introduced in Eq. 7, resulting in a 1.17 dB drop in DPSNR. These experiments highlight that the sole use of 4DGS without the support of our tracking module, does not provide sufficient accuracy for effective motion modeling in autonomous driving scenarios.
Effect of isotropic Gaussian marbles. In Tab. 2(C), we substitute isotropic Gaussian marbles with the original anisotropic Gaussian ellipsoids for dynamic points, which leads to a decrease in DPSNR and DDSIM.
3D tracker / GT pose vs. Our object tracking module. In Tab. 2 (D), we use the 3D object tracker [11, 56] to reconstruct dynamic objects and supervise the motion learning, yielding a 1.3dB decrease in PSNR. This suggests that tracking errors can impact novel view synthesis performance despite our motion learning design, highlighting the necessity of substituting 3D tracker with more generalizable 2D tracker. In Tab. 2 (E), we conduct an experiment using the ground truth object trajectory to benchmark our 2D tracking approach. The relatively small performance gap indicates that our 2D tracking closely approximates the ground truth in our motion learning procedure, supporting robust performance across varied scenarios.
Effect of motion learning module. Compared to Street Gaussians which uses the same tracking results in Tab. 1, the PSNR in Tab. 2 (D) is 0.57 dB higher, indicating the robustness of our motion tracking module to tracking errors.
@@ 5 Conclusion
In this paper, we introduce a novel framework for robust dynamic 3D street scene reconstruction that eliminates the reliance on 3D object trackers. Addressing the generalization limitations of 3D trackers, we propose a robust object tracking strategy based on a 2D foundation model. Our framework also features a motion learning module within an implicit feature space to handle inevitable tracking errors by autonomously refining pose inaccuracies and recovering missed detections. Experiments on the Waymo-NOTR dataset demonstrate its adaptability and superior performance, underscoring its effectiveness in producing high-quality 3D reconstructions across diverse scenarios.

### v2

@@ Abstract
Realistic scene reconstruction in driving scenarios poses significant challenges due to fast-moving objects. Most existing methods rely on labor-intensive manual labeling of object poses to reconstruct dynamic objects in canonical space and move them based on these poses during rendering. While some approaches attempt to use 3D object trackers to replace manual annotations, the limited generalization of 3D trackers – caused by the scarcity of large-scale 3D datasets – results in inferior reconstructions in real-world settings. In contrast, 2D foundation models demonstrate strong generalization capabilities. To eliminate the reliance on 3D trackers and enhance robustness across diverse environments, we propose a stable object tracking module by leveraging associations from 2D deep trackers within a 3D object fusion strategy. We address inevitable tracking errors by further introducing a motion learning strategy in an implicit feature space that autonomously corrects trajectory errors and recovers missed detections. Experimental results on Waymo-NOTR and KITTI show that our method outperforms existing approaches. Our code will be made publicly available.
Figure 1: Comparison of 3D tracker-based Street Gaussians [78] (left) and our approach (right). Existing methods heavily rely on object poses, but 3D trackers struggle with limited generalization [63, 86, 20], leading to flaws in novel view synthesis. In contrast, 2D foundation models show better generalization [77, 73]. Our approach leverages a 2D foundation model [73] for object tracking and learns point motion within an implicit feature space to autonomously correct tracking errors, improving robustness across diverse scenes.
@@ 1 Introduction
Although 2D tracking models demonstrate stronger generalization ability, tracking errors still can exist, especially under adverse conditions such as severe occlusions or distant objects. Since moving the object points solely based on the tracked object pose directly exposes any tracking errors, we aim to go beyond a straightforward reliance on tracked trajectories. Instead, we propose to learn point motion from the predicted trajectory, equipping the model to autonomously identify and correct tracking inaccuracies, recover missed detections, and infer motion in new time steps. An implicit representation is essential for this purpose: 1) It enables the model to refine trajectories without being constrained by bounding box tracks, facilitating smoother and more continuous motion that can adaptively respond to changes. 2) An implicit feature space offers versatility to move each point of an object in a different way, recognizing that vehicles are not strictly rigid objects — for example, doors can open or close. This approach makes it possible to capture subtle, dynamic changes within objects, ultimately producing more accurate reconstructions and enhancing the robustness of novel view synthesis in challenging scenarios.
To this end, we leverage HexPlane representation [7] following 4DGS [70]. HexPlane stores motion-related features by decomposing the 4D spatial-temporal space into six 2D learnable feature planes. A decoder then utilizes these features to predict deformation offsets, dynamically adjusting 3D Gaussians over time. However, 4DGS relies solely on image reconstruction loss for supervision, which is insufficient for handling rapid object motion. When a Gaussian’s initial position is far from its current location, its projection on the image may fall outside the object’s actual area, preventing gradient propagation and hindering optimization. This issue is evident in S3Gaussian [30], which struggles with moving cars due to the lack of explicit motion supervision. To address this, we introduce a training strategy that supervises learned point motion using the predicted trajectory, providing explicit 3D supervision to guide HexPlane in capturing motion dynamics more accurately.
@@
  • We outperform existing methods on Waymo-NOTR and KITTI datasets without relying on ground truth annotations.
@@ 3.3 Tracking Cars in 3D with Robust 2D Tracker
$\tau_{overlap}=\mathcal{O}\cap\mathbf{T}_{t}^{-1}\mathcal{O}_{t},\quad\mathbf{T}_{t}=\operatorname{ICP}(\mathcal{O}_{t},\mathcal{O}).$ (4)
@@ 3.4 Learning Point Motion
Table 1: Comparison with state-of-the-art methods on Waymo-NOTR dataset. StreetGS represents Street Gaussian [78]. The best and the second best results are denoted by pink and blue.
@@ 4.1 Experimental Setup
Datasets. We leverage Waymo-NOTR dataset [79, 66] and KITTI benchmark [21, 22]. The NeRF On-The-Road (NOTR) dataset, introduced by EmerNeRF [79], is a subset of the Waymo Open dataset [66] and includes diverse challenging driving scenarios, such as high-speed, exposure mismatch, and various weather conditions. We use the dynamic32 subset, consisting of 32 dynamic scenes, for evaluating dynamic reconstruction. Following EmerNeRF’s setup [79], we use three frontal camera images resized to $960\times 640$. For scene reconstruction, all image frames are used for training and evaluation; for novel view synthesis, every 10th time step is excluded for evaluation [79]. For the KITTI dataset, we follow the setup of MARS [74] to use 75% or 50% of the images for training with every 4th or every 2nd frame held out for testing, respectively.
Implementation Details. We use the Adam optimizer [36] with the same learning rate schedule as in 3DGS [34]. For long-sequence scene reconstruction, we follow S3Gaussian [30] and segment the scene into multiple clips. The multi-resolution HexPlane encoder has a base resolution of 64, upsampled by factors of 2 and 4 as in [70], and other hyperparameters match those in 3DGS [34]. As LiDAR points lack data for the sky region, we add a plane of points above the maximum height of the scene to represent it following [35]. All experiments are run on a single NVIDIA RTX 3090 GPU, with training taking about 2 hours per video clip and inference speed at 100 FPS at $960\times 640$ resolution.
Baseline Methods. We compare our approach against state-of-the-art methods, including NeRF-based methods (MARS [74], NSG [56], EmerNeRF [79]) and Gaussian-based methods (3DGS [34], StreetGaussian [78], S3Gaussian [30]). To ensure fair comparisons, we apply LiDAR point cloud initialization and add depth regularization to 3DGS. For the Waymo-NOTR dataset, we borrow results of MARS [74] and EmerNeRF [79] from S3Gaussian [30]. We evaluate the 3D-tracker-based method, Street Gaussians [78], using detection results from VoxelNext [11] and associate the detections with SimpleTrack [57]. Our goal is to assess the generalization ability of 3D-tracker-based methods in scenarios without ground truth tracking labels. To emulate a real-world scenario where a 3D tracker trained on a public benchmark dataset is applied to a novel environment, we use pretrained weights from nuScenes [6]—one of the largest and most diverse 3D datasets—and evaluate the model on the Waymo-NOTR dataset. Since NOTR is part of the training set for the Waymo perception task, using Waymo pretrained weights would introduce domain overlap and bias the evaluation, making the nuScenes weights a fairer choice. For a fair comparison, we use pretrained 2D tracker weights from [73] for our method, which have not been trained on the Waymo or KITTI datasets. Please refer to the supplementary material for discussion of the choice of 3D trackers. For the KITTI dataset, we borrow the results of all other methods from Street Gaussians. Please note that Street Gaussians leverages GT tracking annotations on KITTI dataset.
@@ 4.2 Comparisons with the State-of-the-art
KITTI-75% KITTI-50%
Method Extra Input PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
3DGS [34] N/A 19.19 0.737 0.172 19.23 0.739 0.174
NSG [56] GT pose 21.53 0.673 0.254 21.26 0.659 0.266
MARS [74] GT pose 24.23 0.845 0.160 24.00 0.801 0.164
Street Guassians [78] GT pose 25.79 0.844 0.081 25.52 0.841 0.084
Ours 2D tracker [73] 25.49 0.889 0.063 25.11 0.877 0.067
Table 2: Comparison with state-of-the-art methods on KITTI dataset. The best and the second best results are denoted by pink and blue.
@@
The results on the KITTI dataset are presented in Table 2. Our method outperforms the NeRF-based NSG and MARS across all metrics. In the KITTI-75% setting, our approach achieves a 0.045 higher SSIM and a 0.018 lower LPIPS compared to Street Gaussians, although its PSNR is 0.3 dB lower. A similar trend is observed in the KITTI-50% setting. This lower PSNR is primarily because Street Gaussians leverages labor-intensive ground truth tracking annotations, whereas our method uses a generalized 2D tracker—yet still attains comparable performance.
@@ 4.3 Ablation Studies
Method PSNR SSIM DPSNR DSSIM
@@
D 3D tracker 28.02 0.835 24.61 0.746
E D w/o $\mathcal{L}_{\text{motion}}$ 27.54 0.821 23.53 0.703
F GT pose 28.98 0.873 25.71 0.787
G UNINEXT 28.76 0.864 25.43 0.772
H Ours 28.85 0.867 25.58 0.779
@@
Effect of object tracking module. In Tab. 3 (A), we model dynamic points using 4DGS [70] without performing any form of tracking, and observe significant drops across all metrics. This experiment highlight that the sole use of 4DGS does not provide sufficient accuracy for effective motion modeling in autonomous driving scenarios.
Effect of motion learning strategy. In Tab. 3 (B), we use the reconstruction of the tracking module for initialization but remove the motion loss introduced in Eq. 7, resulting in a 1.04 dB drop in DPSNR. This result highlights the importance of our motion learning strategy.
Effect of isotropic Gaussian marbles. In Tab. 3 (C), we substitute isotropic Gaussian marbles with the original anisotropic Gaussian ellipsoids for dynamic points, which leads to a decrease in DPSNR and DSSIM.
3D tracker / GT pose vs. Our object tracking module. In Tab. 3 (D), we use the 3D tracker [11, 57] to reconstruct dynamic objects and supervise motion learning, resulting in a 0.97 dB DPSNR decrease. Without the motion loss, DPSNR decreases further by 1.08 dB (Tab. 3 (E)), showing that our motion learning strategy can compensate for tracking errors, though errors still impact novel view synthesis, emphasizing the need for a more generalizable 2D tracker. In Tab. 3 (F), we conduct an experiment using the ground truth object trajectory to benchmark our 2D tracking approach. The relatively small performance gap indicates that our 2D tracking closely approximates the ground truth in our motion learning procedure, supporting robust performance across varied scenarios.
Choice of 2D tracker. To demonstrate the robustness of our method with respect to the choice of 2D tracker, we employ UNINEXT [77] for 2D tracking. As shown in Table 3 (G), our method maintains stable performance.
@@ 5 Conclusion
In this paper, we introduce a novel framework for robust dynamic 3D street scene reconstruction that eliminates the reliance on 3D object trackers. Addressing the generalization limitations of 3D trackers, we propose a robust object tracking strategy based on a 2D foundation model. Our framework also features a motion learning module within an implicit feature space to handle inevitable tracking errors by autonomously refining pose inaccuracies and recovering missed detections. Experiments on the Waymo-NOTR and KITTI datasets demonstrate its adaptability and superior performance.